# Massless higher-spin supermultiplets in 5D harmonic superspace

### v1

@@ Appendix A Notation and conventions
where $\varepsilon^{mnrsp}$ is completely antisymmetric tensor with $\varepsilon^{01235}=-1$.
@@
The crucial feature of the harmonic superspace is the existence of invariant analytic subspace $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$ in the full superspace. Here the coordinates $x^{m}_{A}$ and $x^{5}_{A}$ are defined as
@@ Appendix B Component structure
It is convenient to study the component structure of prepotentials in the analytic basis with coordinates $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$ which are related with the central basis $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$ through the identities (195). In this basis, the analyticity constraint is satisfied automatically because the corresponding covariant spinor derivative is short, see Eq. (196a).
@@ B.1 Prepotentials in the Wess-Zumino gauge: bosonic fields
In harmonic superspace, 5D higher spin supermultiplets are embedded into three prepotentials $h^{++m_{1}\ldots m_{s-1}}$, $h^{++5m_{1}\ldots m_{s-2}}$ and $h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}}$ subject to the analyticity constraints (93). These constraints may be explicitly solved in the analytic coordinates:
@@
This reach gauge freedom allows us to gauge away infinite tails of auxiliary fields with $SU(2)_{R}$ indices, leaving only the following bosonic components in $\theta$-decompositions of the prepotentials:
@@
As follows from the constraints (90), all these tensors are symmetric and traceless in Lorentz indices $m_{1},m_{2},\ldots,m_{s-1}$. In addition, the component $P^{\hat{\alpha}\hat{\beta}m_{1}\ldots m_{s-2}}$ is $\Gamma$-traceless,
@@
that follows from Eq. (91).
@@
where all tensors are symmetric and traceless in vector indices, whereas $l^{\hat{\alpha}\hat{\beta}m_{1}\ldots m_{s-2}}$ is $\Gamma$-traceless:
@@
Here the rectangles represent Young’s tableau with prime for the tracelessness constraint. Constraint (208) implies the following relations between these components
@@
Next, we will use the gauge freedom represented by the tensors $l_{1}$, $l_{2}$ and $l_{3}$ to gauge away remaining non-physical degrees of freedom and achieve conventional gauge transformations for the physical ones. We do it in the following steps:
@@
  • Eq. (225) dictates the following gauge transformation law for the tensor $A$:
@@
    Recall that $l_{2}^{n(m_{1}\ldots m_{s-2})}$ is traceless with vanishing totally symmetric part because of the gauge fixing imposed in the previous step. Hence, it may be used to gauge away hook-type mixed symmetry components from the tensor $A$, leaving only the symmetric part in it. The number of independent components in $l_{2}$ is
@@
    independent components in the tensor $A$, which is exactly the number of components in a double-traceless tensor of rank (s1)(s-1). Such double-traceless tensor may be decomposed into a pair of totally symmetric and traceless tensors with (s1)(s-1) and (s3)(s-3) Lorentz indices, respectively. This suggests the following decomposition of the tensor $A^{n\,m_{1}\ldots m_{s-1}}$ provided that its mixed-symmetry components are gauged away using the gauge freedom in $l_{2}$:
@@
    Remembering that $l_{2}^{n(m_{1}\ldots m_{s-2})}$ is traceless, from Eq. (227) we find gauge transformation of the tensor $A_{2}$:
@@
    that accounts for the components of the tensor $l_{2}=\left(\begin{array}[]{|c|}\hline\cr\phantom{m}\\ \hline\cr\end{array}\times\begin{array}[]{|c|}\hline\cr s-2\\ \hline\cr\end{array}\,^{\prime}\right)^{\prime}$. Hence, after using tensors $l_{1}$ and $l_{2}$ for eliminating other components considered above, we are left with
@@
    were $B_{1}$ and $B_{2}$ are totally symmetric and traceless tensors,
@@
    This double-traceless superfield cannot be embedded as component into the superfield prepotential $h_{A}^{++m_{1}\ldots m_{s-1}}$ because the correct component of this superfield is given by Eq. (262).
@@
Table 1: Number of bosonic and fermionic degrees of freedom contained in the prepotentials describing spin-$s$ supermultiplet off shell. All gauge degrees of freedom are eliminated.
Table 1 summarizes the number of bosonic degrees of freedom left in the prepotentials (203) after elimination of all nonphysical and pure gauge degrees of freedom,
@@ B.2 Prepotentials in the Wess-Zumino gauge: fermionic fields
Here $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$, $\chi^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ and $\rho^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ are symmetric and traceless tensors in all vector indices. The latter field obeys also $\Gamma$-tracelessness constraint, $(\Gamma_{m_{1}})_{\hat{\alpha}\hat{\beta}}\rho^{m_{1}\ldots m_{s-2}\hat{\beta}i}=0$, which follows from Eq. (91).
Note that the last two terms in Eq. (294a) may be eliminated by a redefinition of $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$. However, we keep these terms because they help represent gauge variation of $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$ in a suitable form (298) below.
@@
The auxiliary field $\rho^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ and the gauge parameter $\epsilon^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ contain the same number of degrees of freedom since they respect the same symmetry and constraints. Therefore, for counting fermionic degrees of freedom it is sufficient to consider only the fields $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$ and $\chi^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ which are both symmetric and traceless in vector indices, but the $\Gamma$-tracelessness is not imposed. Recalling that a symmetric traceless tensor of rank $s$ contains $\left(\begin{array}[]{c}s+4\\ s\end{array}\right)-\left(\begin{array}[]{c}s+2\\ s-2\end{array}\right)$ independent components, we conclude that the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ contain
@@ B.3 Quadratic action for bosonic components
In addition to the prepotentials $h^{++m(s-1)}$ and $h^{++5m(s-2)}$, the action (153) involves also the superfields $h^{--m(s-1)}$ and $h^{--5m(s-2)}$, which are related to the former ones by the zero-curvature equations (101). Substituting the expressions (293) into the zero-curvature equations (101) it is possible to find the component structure of the superfields $h^{--m(s-1)}$ and $h^{--5m(s-2)}$. Then, using the following conventions for Grassmann and harmonic integrations
@@
The auxiliary fields $V^{ij\,m(s-1)}$ and $D^{ij\,m(s-2)}$ have trivial equations of motion and completely decouple. The auxiliary field $P^{\hat{\alpha}\hat{\beta}\,m(s-2)}$, however, couples non-trivially with the physical components. This situation is common for supersymmetric higher spin field theories, see, e.g., KuzenkoSibiryakov94 . In the present case, elimination of this auxiliary field $P^{\hat{\alpha}\hat{\beta}\,m(s-2)}$ represents a non-trivial task because this field is $\Gamma$-traceless, see Eq. (204). As a consequence, the components of this field
@@
For the spin-2 supermultiplet, the $\Gamma$-tracelessness (204) is not required, and the constraints (313) are not imposed. In this case, the action (311) simplifies:
@@
Eliminating the remaining auxiliary fields,
@@
we obtain the quadratic action for physical fields of spins 1 and 2:

### v2

@@ Appendix A Notation and conventions
where $\varepsilon^{mnrsp}$ is the completely antisymmetric tensor with $\varepsilon^{01235}=-1$.
@@
The crucial feature of the harmonic superspace is the existence of an invariant analytic subspace $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$ in the full superspace. Here the coordinates $x^{m}_{A}$ and $x^{5}_{A}$ are defined as
@@ Appendix B Component structure
It is convenient to study the component structure of prepotentials in the analytic basis with coordinates $(x^{m}_{A},\theta^{+}_{\hat{\alpha}},u^{\pm}_{i},x^{5}_{A})$ which are related to the central basis $(x^{m},\theta^{\pm}_{\hat{\alpha}},u^{\pm}_{i},x^{5})$ through the identities (200). In this basis, the analyticity constraint is satisfied automatically because the corresponding covariant spinor derivative is short, see Eq. (201a).
@@ B.1 Prepotentials in the Wess-Zumino gauge: bosonic fields
In harmonic superspace, 5D higher spin supermultiplets are embedded into three prepotentials $h^{++m_{1}\ldots m_{s-1}}$, $h^{++5m_{1}\ldots m_{s-2}}$ and $h^{++\hat{\alpha}+m_{1}\ldots m_{s-2}}$ subject to the analyticity constraints (98). These constraints may be solved explicitly in the analytic coordinates:
@@
This gauge freedom allows us to gauge away infinite tails of auxiliary fields with $SU(2)_{R}$ indices, leaving only the following bosonic components in $\theta$-decompositions of the prepotentials:
@@
As follows from the constraints (95), all of these tensors are symmetric and traceless in Lorentz indices $m_{1},m_{2},\ldots,m_{s-1}$. In addition, the component $P^{\hat{\alpha}\hat{\beta}m_{1}\ldots m_{s-2}}$ is $\Gamma$-traceless,
@@
which follows from Eq. (96).
@@
where all tensors are symmetric and traceless in vector indices, and $l^{\hat{\alpha}\hat{\beta}m_{1}\ldots m_{s-2}}$ is $\Gamma$-traceless:
@@
Here the rectangles represent Young’s tableau with a prime for the tracelessness constraint. The constraint (213) implies the following relations between these components
@@
Next, we will use the gauge freedom represented by the tensors $l_{1}$, $l_{2}$ and $l_{3}$ to gauge away any remaining non-physical degrees of freedom and achieve the conventional gauge transformations for the physical ones. We do this in the following steps:
@@
  • Eq. (230) gives the following gauge transformation law for the tensor $A$:
@@
    Recall that $l_{2}^{n(m_{1}\ldots m_{s-2})}$ is traceless with vanishing totally symmetric part because of the gauge fixing imposed in the previous step. Hence, it may be used to gauge away hook-type mixed symmetry components from the tensor $A$, leaving only the symmetric part. The number of independent components in $l_{2}$ is
@@
    independent components in the tensor $A$, which is exactly the number of components in a double-traceless tensor of rank (s1)(s-1). Such a double-traceless tensor may be decomposed into a pair of totally symmetric and traceless tensors with (s1)(s-1) and (s3)(s-3) Lorentz indices, respectively. This suggests the following decomposition of the tensor $A^{n\,m_{1}\ldots m_{s-1}}$ provided that its mixed-symmetry components are gauged away using the gauge freedom in $l_{2}$:
@@
    Remembering that $l_{2}^{n(m_{1}\ldots m_{s-2})}$ is traceless, from Eq. (232) we find the gauge transformation of the tensor $A_{2}$:
@@
    which accounts for the components of the tensor $l_{2}=\left(\begin{array}[]{|c|}\hline\cr\phantom{m}\\ \hline\cr\end{array}\times\begin{array}[]{|c|}\hline\cr s-2\\ \hline\cr\end{array}\,^{\prime}\right)^{\prime}$. Hence, after using tensors $l_{1}$ and $l_{2}$ for eliminating other components considered above, we are left with
@@
    where $B_{1}$ and $B_{2}$ are totally symmetric and traceless tensors,
@@
    This double-traceless superfield cannot be embedded as a component into the superfield prepotential $h_{A}^{++m_{1}\ldots m_{s-1}}$ because the correct component of this superfield is given by Eq. (267).
@@
Table 1: Number of bosonic and fermionic degrees of freedom contained in the prepotentials describing a spin-$s$ supermultiplet off shell. All gauge degrees of freedom are eliminated.
Table 1 summarizes the number of bosonic degrees of freedom left in the prepotentials (208) after elimination of all unphysical and pure gauge degrees of freedom,
@@ B.2 Prepotentials in the Wess-Zumino gauge: fermionic fields
Here $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$, $\chi^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ and $\rho^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ are symmetric and traceless tensors in all vector indices. The latter field also obeys the $\Gamma$-tracelessness constraint, $(\Gamma_{m_{1}})_{\hat{\alpha}\hat{\beta}}\rho^{m_{1}\ldots m_{s-2}\hat{\beta}i}=0$, which follows from Eq. (96).
Note that the last two terms in Eq. (299a) may be eliminated by a redefinition of $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$. However, we keep these terms because they help represent the gauge variation of $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$ in a suitable form (303) below.
@@
The auxiliary field $\rho^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ and the gauge parameter $\epsilon^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ contain the same number of degrees of freedom since they respect the same symmetry and constraints. Therefore, when counting the fermionic degrees of freedom it is sufficient to consider only the fields $\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}$ and $\chi^{m_{1}\ldots m_{s-2}\hat{\alpha}i}$ which are both symmetric and traceless in vector indices, but the $\Gamma$-tracelessness is not imposed. Recalling that a symmetric traceless tensor of rank-$s$ contains $\left(\begin{array}[]{c}s+4\\ s\end{array}\right)-\left(\begin{array}[]{c}s+2\\ s-2\end{array}\right)$ independent components, we conclude that the prepotentials $h^{++Mm_{1}\ldots m_{s-2}}$ contain
@@ B.3 Quadratic action for bosonic components
In addition to the prepotentials $h^{++m(s-1)}$ and $h^{++5m(s-2)}$, the action (158) also involves the superfields $h^{--m(s-1)}$ and $h^{--5m(s-2)}$, which are related to the former ones by the zero-curvature equations (106). By substituting the expressions (298) into the zero-curvature equations (106) it is possible to find the component structure of the superfields $h^{--m(s-1)}$ and $h^{--5m(s-2)}$. Then, using the following conventions for Grassmann and harmonic integrations
@@
The auxiliary fields $V^{ij\,m(s-1)}$ and $D^{ij\,m(s-2)}$ have trivial equations of motion and completely decouple. The auxiliary field $P^{\hat{\alpha}\hat{\beta}\,m(s-2)}$, however, couples non-trivially with the physical components. This situation is common for supersymmetric higher spin field theories, see, e.g., KuzenkoSibiryakov94 . In the present case, elimination of this auxiliary field $P^{\hat{\alpha}\hat{\beta}\,m(s-2)}$ is a non-trivial task because this field is $\Gamma$-traceless, see Eq. (209). As a consequence, the components of this field
@@
For the spin-2 supermultiplet, the $\Gamma$-tracelessness (209) is not required, and the constraints (318) are not imposed. In this case, the action (316) simplifies to
@@
By eliminating the remaining auxiliary fields,
@@
we obtain the quadratic action for the physical fields of spins 1 and 2:
@@
B.4 Quadratic action for fermionic components
Recall that the fermionic contributions to the prepotentials
$h_{A}^{++m(s-1)}\,,\quad h_{A}^{++5m(s-2)}\,,\quad h^{++\hat{\alpha}+m(s-2)}$ (324)
are given in Eqs. (299a), (299b) and (299c), respectively. Let us denote
$\tilde{\psi}^{m_{1}\ldots m_{s-1}\hat{\alpha}i}=$ $\,\psi^{m_{1}\ldots m_{s-1}\hat{\alpha}i}+\frac{1}{2}(\Gamma^{(m_{1}})^{\hat{\alpha}}{}_{\hat{\beta}}\chi^{m_{2}\ldots m_{s-1})\hat{\beta}i}$
$-\frac{s-2}{2s-1}\eta^{(m_{1}m_{2}}(\Gamma_{n})^{\hat{\alpha}}{}_{\hat{\beta}}\chi^{m_{3}\ldots m_{s-1})n\hat{\beta}i}\,$. (325)
Then $h_{A}^{++m(s-1)}$ is given by
$h_{A}^{++m(s-1)}=-2(\theta^{+})^{2}\theta_{\hat{\alpha}}^{+}\tilde{\psi}^{m(s-1)\hat{\alpha}i}u^{-}_{i}\,.$ (326)
Just like in the bosonic case, one can use the zero-curvature equations (106) to find the fermionic component structure of the superfields $h^{--m(s-1)}$, $h^{--5m(s-2)}$. Performing Grassmann integration using Eq. (314) and harmonic integration using
$\int\mathrm{d}u\,f^{i}u^{+}_{i}g^{j}u^{-}_{j}=\frac{1}{2}f^{i}g^{j}\varepsilon_{ij}=\frac{1}{2}f^{i}g_{i}\,$, (327)
we obtain the action for the fermionic components
$S_{\mathrm{ferm}}=$ $\,S_{1}-\frac{2s-1}{s-1}S_{2}\,$, (328)
$S_{1}=$ $\,\frac{\mathrm{i}}{8}\int{\mathrm{d}}^{5}x\bigg\{\tilde{\psi}^{m(s-1)\hat{\alpha}i}(\Gamma^{n})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\tilde{\psi}_{m(s-1)\hat{\beta}i}-\tilde{\psi}^{m_{1}\dots m_{s-1}\hat{\alpha}i}(\Gamma_{m_{1}})_{\hat{\alpha}}{}^{\hat{\beta}}\rho_{m_{2}\dots m_{s-1}\hat{\beta}i}\bigg\}\,$,
$S_{2}=$ $\,\frac{\mathrm{i}}{32}\int{\mathrm{d}}^{5}x\bigg\{\chi^{m(s-2)\hat{\alpha}i}(\Gamma^{n})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\chi_{m(s-2)\hat{\beta}i}-2\chi^{m(s-2)\hat{\alpha}i}\rho_{m(s-2)\hat{\alpha}i}\bigg\}\,.$
The component fields $\rho^{m(s-2)\hat{\alpha}i}$ and the $\Gamma$-traceless part of $\chi^{m(s-2)\hat{\alpha}i}$, denoted by $\chi^{\prime m(s-2)\hat{\alpha}i}$ in Section B.2, are auxiliary. The component fields $\psi^{m(s-1)\hat{\alpha}i}$ and the $\Gamma$-trace of $\chi^{m(s-2)\hat{\alpha}i}$, denoted by $\phi^{m(s-3)\hat{\alpha}i}$ in Section B.2, combine into the Fang–Fronsdal field (312).
Let us consider the case $s=2$ in more detail. In this case $\tilde{\psi}^{m\hat{\alpha}i}$ is given by
$\tilde{\psi}^{m\hat{\alpha}i}=\psi^{m\hat{\alpha}i}+\frac{1}{2}(\Gamma^{m})^{\hat{\alpha}}{}_{\hat{\beta}}\chi^{\hat{\beta}i}\,$, (329)
and the action simplifies to yield
$S_{\mathrm{ferm}}=$ $\int{\mathrm{d}}^{5}x\bigg\{\frac{\mathrm{i}}{8}\psi^{m\hat{\alpha}i}(\Gamma^{n})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\psi_{m\hat{\beta}i}+\frac{\mathrm{i}}{8}\chi_{\hat{\beta}i}\partial_{n}\psi^{n\hat{\beta}i}+\frac{\mathrm{i}}{4}(\Sigma^{mn})_{\hat{\alpha}}{}^{\hat{\beta}}\chi^{\hat{\alpha}i}\partial_{n}\psi_{m\hat{\beta}i}$
$-\frac{\mathrm{i}}{8}\psi^{m\hat{\alpha}i}(\Gamma_{m})_{\hat{\alpha}}{}^{\hat{\beta}}\rho_{\hat{\beta}i}-\frac{\mathrm{i}}{8}\chi^{\hat{\alpha}i}\rho_{\hat{\alpha}i}\bigg\}\,.$ (330)
From the equations of motion for $\rho_{\hat{\beta}i}$ and $\chi_{\hat{\beta}i}$ we find that these fields are solved in terms $\psi^{m\hat{\alpha}i}$ as follows
$\chi^{\hat{\alpha}i}=$ $\,(\Gamma_{m})^{\hat{\alpha}}{}_{\hat{\beta}}\psi^{m\hat{\beta}i}\,$, (331)
$\rho_{\hat{\alpha}i}=$ $\,\partial^{n}\psi_{n\hat{\alpha}i}+2(\Sigma^{mn})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\psi_{m\hat{\beta}i}\,.$ (332)
Substituting these expressions back into the action (330) gives
$S_{\mathrm{ferm}}=-\frac{\mathrm{i}}{8}\int{\mathrm{d}}^{5}x\,\psi^{m\hat{\alpha}i}\Big[2(\Gamma_{m}\Sigma^{pn})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\psi_{p\hat{\beta}i}-(\Gamma^{n})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\psi_{m\hat{\beta}i}+(\Gamma^{n})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{m}\psi_{n\hat{\beta}i}\Big]\,.$ (333)
One can show that Eq. (333) can also be presented in the form
$S_{\mathrm{ferm}}=-\frac{\mathrm{i}}{8}\int{\mathrm{d}}^{5}x\,\psi^{m\hat{\alpha}i}\eta_{ms}(\Gamma^{[s}\Gamma^{n}\Gamma^{p]})_{\hat{\alpha}}{}^{\hat{\beta}}\partial_{n}\psi_{p\hat{\beta}i}\,$, (334)
which is the standard action for a spin-$3/2$ field.